Homestake Meeting Minutes, 11/13/14
Attending: Daniel, Vuk, Riccardo, Tanner, Pat, Nelson
Homestake wiki: https://zzz.physics.umn.edu/groups/homestake/home
1) Homestake: planning the next trip
Vuk: PASSCAL shipment should be there, talked with Bryce. Purchased 5 sealed batteries for the surface stations, and two 12V 1.25 Amp batteries with smart chargers for two underground stations. The third underground station should be powered with a DC power supply. Tanner and Pat will bring these from MSP, pick up Gary from Rapid airport, and pick up the doghouses from the Tractor Supply store in Rapid.
On Monday will have to open crates at the Ross Warehouse and check the equipment. Also install the radio on the roof of Yates admin building and the GPS transceiver in the IT room. Will need Jamey for this.
On Tuesday, Gary and Pat can start installing surface stations, and Tanner and Jamey can install the three stations at 4850. The remaining days will be used to complete the surface stations and set up the data flow.
2) UMN test setup
Tanner: Trying to find a good spot to test the GPS transceivers, but could not do it. Today will try to do it at the roof of our physics building, to confirm that transceivers work.
Have been working with PASSCAL to get the configuration files. Ran into a problem with the laptop, should know more over the next day or two.
Building cables for GPS-Q330 connection, should receive the necessary connectors on Monday/Tuesday at Homestake. Will have to build the cables at Homestake on Monday or Tuesday evening.
Vuk: this may require us to rearrange the schedule. Maybe use Tue to set up the first station at 4850 (Davis), and to set up the software and data flow on the surface. In this case, as Pat and Gary make new surface stations, they could be added to the data pipeline.
Vuk: Should get access to the existing computer on the surface. Tanner will contact Shivaraj. Should decide how to move the data around, the existing surface machine should be useful.
3) Analysis of existing data
Vuk, Noah, and Eric met last week, believe we found the reason why some of the R-wave recovery did not work. Noah will test the correction over the coming week.